There are 10 stalls for animals in an exhibition. Three animals; lion, pussycat




and horse are to be exhibited. Animals of each kind are not less than 10 in




number. What is the possible number of ways of arranging the exhibition?

1
Expert's answer
2021-12-27T17:56:42-0500

For each stall there are 3 ways to exhibit one of three animals. Since there are 10 stalls, by Combinatorial Product Rule, the number of ways of arranging the exhibition is equal to "3\cdot 3\cdot\ldots\cdot 3=3^{10}=59,049."